OpenStudy (anonymous):

Solve the equation 2sin^2θ -sinθ -1=0

OpenStudy (jdoe0001):

notice is just a quadratic equation

OpenStudy (dls):

Let sinθ=t and solve.

OpenStudy (akashdeepdeb):

Let sin θ = x 2x2 - x - 1 = 0 Solve this quadratic equation to get 2 values of x and sinθ Understood? :)

OpenStudy (jdoe0001):

\(\bf 2sin^2(\theta) - sin(\theta) - 1=0 \quad ,\color{blue}{sin(\theta) = t},\quad \implies 2t^2-t-1=0\)

OpenStudy (anonymous):

so would the answer just be -1/2 and 1?

OpenStudy (akashdeepdeb):

Yes. :) Find θ just like that.